OpenStudy (anonymous):

medal for answer. -4 , 8 , -16 , 32 . . . a. arithmetic; 64, 128, 256 b. geometric, -64 , 128 , -256 c. geometric, -48 64 , -80 d. neither

OpenStudy (cwrw238):

-4,8,-16,32 is a geometric sequence - to get next term you multiply by -2

OpenStudy (cwrw238):

the next 3 terms are -64,128,-256

OpenStudy (cwrw238):

b is correct

OpenStudy (anonymous):

can you help me with 2 more please?

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

81, 27, 9, 3 . . . .

OpenStudy (anonymous):

a. arithmetic , 0, -3 , -6 b. geometric, 0, -3 , -6 c. geometric, 1, 1/3, 1/9 d. neither.

OpenStudy (cwrw238):

hint: divide 27 by 81 9 by 27 3 by 9 what do you get?

OpenStudy (cwrw238):

if you get the same value for all 3 divisions then you have a geometric sequence

OpenStudy (anonymous):

i did

OpenStudy (cwrw238):

what was the value?

OpenStudy (cwrw238):

it should be 1/3 right?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

ok so now the next 3 terms are 3 * 1/3 = 1 1* 1/3 = 1/3 and 1/3 * 1/3 = 1/9 do you follow that?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

so thats option c sorry but i must go